OpenStudy (anonymous):

Write the slope intercept equation for the line with slope -2/3 and the y intercept (0,3)

Parth (parthkohli):

The form of an equation in the slope intercept form is y = mx + b...m is the slope and b is the y-intercept.

Parth (parthkohli):

y = -2/3x + b 3 = b (x = 0 so the fraction = 0) \(\Huge \color{MidnightBlue}{\Rightarrow y = {-2 \over 3}x + 3 }\)